A company manufactures two products, A and B. Ech unit of A requires 3 labor hours and each unit of requires 5 labor hours. Dally manufacturing capacity 150 labor hours.
1) If x units of Product A and y units of product are manufactured each and all labor hours are to
used, determine the linear equation that requires the use of 150 labor hours per day.
ii) How many units of A can be made each day if 2units of B are manufactured each day?
iii) How many units of A can be made each week if 25 units of B are manufactured each day? (Assume a day work week)



Answer by greenestamps(13200)



(i) ANSWER: 3x+5y=150

x and y are non-negative integers.

Since 5y is a multiple of 5 and the total 150 is a multiple of 5, 3x must also be a multiple of 5; and that means s must be a multiple of 5.

(ii) If 2 units of B are manufactured each day, then y=2, and

3x+10=150
3x=140

But 140 is not a multiple of 3 -- so 2 units of B manufactured each day violates the requirement that all 150 labor hours be used.

ANSWER: There is no solution to this problem

(iii) This problem talks about production in a week, assuming a 5-day work week. But we have the same problem here as in (ii) -- if y=25 then daily production is

3x+125=150
3x=25

And again 25 is not a multiple of 3, so the requirement of using all 150 labor hours can't be met.

ANSWER: Again there is no solution to this problem.